### v2

@@ Abstract
Experiments deliver credible treatment effects estimates, but are often localized to specific sites, populations, or mechanisms. When such estimates are insufficient to extrapolate effects for broader policy questions as for external validity and general-equilibrium (GE) effects, researchers combine trials with external evidence from reduced-form or structural observational estimates, or prior experiments. We develop a unified framework for designing experiments in this setting: the researcher selects which parameters (or moments) to identify experimentally from a feasible set of possible experiments, allocates sample size, and specifies how to weight experimental and observational estimators. Because observational inputs may be biased in ways unknown ex ante, we adopt a minimax proportional-regret objective that evaluates any candidate design relative to an oracle that knows the bias and jointly chooses the design and estimator. This yields a transparent bias-variance trade-off that requires no prespecified bias bound and depends only on information about the precision of the estimators and the estimand’s sensitivity to underlying parameters. We illustrate the framework by (i) designing small-scale cash-transfer experiments aimed at estimating GE effects and (ii) optimizing site selection for microfinance interventions.
@@ 1 Introduction
Randomized controlled trials (RCTs) have improved empirical economics by providing internally valid causal estimates. However, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism. These effects, although useful, are often not sufficient to answer broader questions about external validity and (general) equilibrium effects of at-scale interventions. For instance, the median evaluation of experiments in top economics journals is “representative of a population of 10,885 units, studied a treatment delivered to 5,340 units, and randomized in clusters of 26 units per cluster” (Muralidharan and Niehaus, 2017).
@@
In response, a growing literature in development economics (e.g., Meghir et al., 2022; Bassi et al., 2022; Attanasio et al., 2012; de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), labor economics (e.g., Chetty et al., 2016), and industrial organization (e.g., Katz and Allcott, 2025; Allcott et al., 2025) combine experiments with external evidence—observational estimates and/or results from experiments in other contexts—to estimate complex counterfactuals that no single experiment can identify. In our review of AEA journals over the past decade, more than 30% of experimental papers complement RCT evidence with at least one observational input, either reduced-form or structural (Figure 12). This trend poses a design question: given constraints on feasible experiments, which experiments should be run (and how) to estimate such counterfactuals?
@@
Our main contribution is a framework for designing experiments combined with external evidence. By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other contexts. We provide a procedure that selects which experiments to run subject to a budget (e.g., which treatment arms and/or sub-populations to include), sets their precision via sample allocation, and prescribes how to combine experimental and external evidence. External input may be biased in ways not known ex ante (e.g., due to confounding or lack of external validity).
@@
For an illustrative example, consider a government piloting a cash-transfer program in a small set of districts to increase children’s school attendance. The trial delivers a local average effect, but policy decisions often require counterfactuals at scale, allowing prices and wages to adjust (e.g. Todd and Wolpin, 2006; Egger et al., 2022). Researchers therefore combine experimental evidence with a supply-demand model that leverages observational data to map local impacts into economy-wide outcomes. The design question is twofold: (i) which experiment to run (i.e., which parameters/effects are the most valuable to learn experimentally) given feasibility constraints and (ii) how to allocate sample across sites and arms so that, once combined with external inputs, we precisely estimate the effect at scale.
@@
The first step is to specify the estimand and the underlying parameters. Define $\tau(\theta)$ the (policy-relevant) object of interest, where $\tau$ is a known smooth function and $\theta$ is a vector of unknown parameters; $\tau$ may be scalar or vector-valued. Researchers also observe external baseline estimates of $\theta$—henceforth observational estimates—which may be biased. Each feasible experiment can identify a subset of parameters without bias. This parameterization makes explicit which components of $\theta$ are learned experimentally and which may rely on biased observational evidence; not all components need (or can) be learned experimentally.
@@
Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in rural Kenya receive the subsidy. The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium , and (iii) wage adjustments that shift the returns to schooling. Due to cost constraints, researchers may only run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$. They may choose between a CCT treatment arm to estimate the direct effect, or an unconditional-cash arm (UCT) to measure the income elasticity, and extrapolate the GE effects using evidence from a previous study in Mexico (Todd and Wolpin, 2006). (When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ targets the average effect across sites.)
The second step is to design the experiment: (i) selecting which experiment to run; (ii) allocating sample sizes across arms; and (iii) choosing how to combine observational and experimental evidence. We allow flexibility in (iii): researchers may choose to either optimally combine both sources when available or rely solely on experimental estimates. Parameters not identified experimentally are based on observational estimates.
@@
A natural starting point is to design the experiment to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps worst-case over the observational bias. In practice, however, the size and even the direction of bias are unknown (before the experiment). A worst-case approach would therefore be overly-conservative and disregard information about the variance. We instead adapt the definition of proportional (adaptation) regret previously studied when combining two estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem (with multiple estimators). Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the estimator. Taking its supremum over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
@@
Our main result gives an explicit novel characterization of proportional regret. For any design and weights that combine the experimental and observational point estimates, the regret is the maximum of two normalized components. The first is a variance regret: the estimator’s sampling variance under the chosen design and weights, divided by the smallest attainable variance. The second is a bias regret: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias over the class of designs. The variance regret depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters. This gradient is evaluated, under mild conditions, at the observational estimates. [^4] For this result to hold, we assume that $\tau(\theta)$ is twice differentiable with bounded derivatives. For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al. (2020); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error. Section 5.1 provides details. The bias regret depends only on the sensitivity vector and the shrinkage weights, known ex-ante.
@@
This characterization makes the bias-variance trade-off transparent. At the optimum, the two normalized components tend to be equalized: when the bias component dominates, the design prioritizes high-sensitivity parameters; when variance dominates, the procedure invests sample where it most reduces variance and relies more on the most precise observational inputs. This yields a nested procedure that chooses jointly the estimator, the sample allocation and the experiment to run, trading off precision with sensitivity of the policy estimand. The resulting design and estimator only require knowledge of the expected variance-covariance matrix (standard in experimental design problems, Gerber and Green, 2012) and of the observational estimates. It can then be reported in a pre-analysis plan. That is, the inputs for the design are the same that would be required by standard power-analysis plans.
From a theoretical perspective, our result leverages a key insight. Because here bias arise from multiple parameters, the choice of the uncertainty set is pivotal. While we do not require researchers to specify such a set, regret is with respect to an oracle that knows the worst-case bias within an $\ell_{\infty}$-ball. Compared with $\ell_{1}$ or $\ell_{2}$, the $\ell_{\infty}$-ball prevents that a large bias in one component is “offset” by smaller biases elsewhere. This is attractive when biases may be positively correlated across parameters. We show that the $\ell_{\infty}$ choice combined with our design problem (with decisions made before experimental data is realized), yields a quasi-convex regret in the bias radius as the oracle optimizes over the designs. This result, worst-case over the radius, delivers a transparent and easy-to-compute variance-bias trade-off.
We then extend the model to general method-of-moments estimators that combine experimentally identified moments with observational moments. As in the baseline case, observational moments may incur unknown misspecification. We optimize (i) which moments to identify experimentally and (ii) how to combine experimental and observational moments to estimate the counterfactual of interest. Our main result continues to hold in this more general scenario provided that the moment estimators are (asymptotically) linear in the moments. Linearity occurs (under smoothness conditions) as long as the ratio of the bias to standard error is no faster than $n^{1/4}$, with $n$ denoting the observational sample size; standard local-asymptotics with the bias and standard error of the same order is a special case.
We illustrate the framework in the cash‐transfer application, when the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006; Attanasio et al., 2012). We are concerned that these preliminary estimates may lack external validity in Kenya. We compare three designs: (i) a CCT arm to estimate the direct schooling effect; (ii) an unconditional cash transfer (UCT) arm to estimate income effects; and (iii) a two-arm design that allocates sample across CCT and UCT under a common budget. Running a single arm captures settings with fixed costs per arm; allowing both arms with budgeted allocation captures cases with non-binding fixed costs but binding variable costs. When both arms are feasible, the regret-optimal allocation puts most participants in the CCT and a small fraction in the UCT, reflecting the target’s greater sensitivity to the direct effect despite the CCT’s higher variance. If only one arm can be run, the choice hinges on sample size: for very small $n$, the lower-variance of the UCT arm makes this preferred over a (too noisy) CCT arm; once $n\geq 500$, the more informative but noisier CCT yields lower regret. Relative to an oracle that knows the bias, the optimal design closely tracks the oracle. We provides a step-by-step guide for practitioners applicable in other contexts.
@@
As a second application, we design where to run a microfinance experiment in rural India, integrating evidence from earlier nonrandomized microfinance introductions. We start from the observational estimates in Banerjee et al. (2024) and design an experiment that selects one or more areas for randomization. Because Banerjee et al. (2024) also implement a separate randomized expansion, we can calibrate performance of our design (that is agnostic of the bias) by comparing each candidate design’s MSE to that of an oracle that knows the bias, using the experimental estimates to calibrate the bias. The optimal design concentrates sample in high-sensitivity, high-variance areas and then spreads as the budget grows; compared to a benchmark that randomly chooses areas and splits sample equally, it reduces MSE by roughly 20% under the calibrated bias.
@@
In summary, we complement this literature by studying which experiment to run (and how) when not all parameters relevant for the target estimand can be learned experimentally.
Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design. References include Manski and Tetenov (2016), Banerjee et al. (2020), Manski (2004), Dominitz and F. Manski (2017), Olea et al. (2024), Hu et al. (2024), Breza et al. (2025) among others. These references focus on settings where researchers have only access to experimental variation, instead of combining it with observational evidence, motivating different designs and objective functions. In the context of site-selection, we complement literature that leverages correctly-specified models (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2025; Andrews et al., 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024). We adapt the worst-case proportional regret idea from Armstrong et al. (2024) and Tsybakov (1998) here studied in the context of experimental design. Our definition of sensitivity of the estimand to each parameter builds on sensitivity analysis in Andrews et al. (2020). Unlike both strands of this literature where the design is fixed and researchers optimize over the choice of the estimator only, here we optimize the design itself. The objective and analysis change: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
2 Problem description
Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional unknown parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$. For simplicity, we assume that the estimand is univariate, while Section 5.3 shows how our framework directly extends to multivariate $\tau(\theta)\in\mathbb{R}^{q}$ for $q>1$.
The goal is to construct an estimator $\hat{\tau}$ that accurately approximates $\tau(\theta)$ by combining existing evidence (e.g., observational studies) with experimental variation designed by the researcher. Our question is how to design such experiments under feasibility constraints.

### v3

@@ Abstract
Experiments deliver credible treatment-effect estimates but, because they are costly, are often restricted to specific sites, small populations, or particular mechanisms. A common practice across several fields is therefore to combine experimental estimates with reduced-form or structural external (observational) evidence to answer broader policy questions such as those involving general equilibrium effects or external validity. We develop a unified framework for the design of experiments when combined with external evidence, i.e., choosing which experiment(s) to run and how to allocate sample size under arbitrary budget constraints. Because observational evidence may suffer bias unknown ex-ante, we evaluate designs using a minimax proportional-regret criterion that compares any candidate design to an oracle that knows the observational study bias and jointly chooses the design and estimator. This yields a transparent bias-variance trade-off that does not require the researcher to specify a bias bound and relies only on information already needed for conventional power calculations. We illustrate the framework by (i) designing cash-transfer experiments aimed at estimating general equilibrium effects and (ii) optimizing site selection for microfinance interventions.
@@ 1 Introduction
Randomized controlled trials (RCTs) have improved empirical economics by providing internally valid estimates. However, in practice, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism, as Muralidharan and Niehaus (2017) document. [^4] Experiments in top economics journals are in median “representative of a population of 10,885 units […] and randomized in clusters of 26 units per cluster” (Muralidharan and Niehaus, 2017).These effects, although useful, are often not sufficient to answer relevant policy questions about external validity and general equilibrium (GE) effects of at-scale interventions.
@@
In response, a large literature in development (e.g., Meghir et al., 2022; Todd and Wolpin, 2006; Attanasio et al., 2012; de Albuquerque et al., 2025), education and labor economics (e.g., Allende et al., 2019; Larroucau et al., 2024; Chetty et al., 2016), and more recently industrial organization (e.g., Katz and Allcott, 2025; Allcott et al., 2025) combines experiments with external evidence—observational and/or results from experiments in other settings—to extrapolate policy-relevant counterfactuals that no single experiment can identify. In our review of AEA journals over the past decade, about one-third of experimental papers complement experimental with observational inputs, either reduced-form or structural (Figure 11). Since budget constraints often bind, such common practice raises the question of how to choose the most informative experiment(s) and design for policy-relevant estimands from a (typical large) menu of possible experiments (e.g. Ravallion, 2012; Niederle, 2025).
@@
Our main contribution is a framework for choosing which experiment to run and how to design it in this common setting of combining experimental with external evidence. By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments in other contexts. External evidence may be biased in ways not known ex ante (e.g., due to confounding or lack of external validity). We provide a procedure that (1) helps researchers disentangle the role of each experiment in this setting, (2) selects which experiments to run subject to a budget (e.g., which treatment arms and/or sub-populations), (3) optimally allocates sample size, and (4) prescribes how to combine experimental and external evidence for estimation. We allow for arbitrary (user-specific) constraints in each step, for both the choice of the design and of the estimator. [^5] For (4), researchers may constrain the estimator to either only use experimental evidence when available or optimally combine experimental with (biased) observational evidence for estimating certain parameters.
@@
For an illustrative example, consider a government piloting a cash-transfer program in a small set of districts to increase children’s school attendance. The trial delivers a local average effect, but policy decisions often require counterfactuals at scale, allowing prices and wages to adjust (e.g. Todd and Wolpin, 2006; Egger et al., 2022). Researchers therefore combine experimental evidence with a supply-demand model that leverages observational data to map local impacts into economy-wide outcomes. The design question is (i) which mechanisms are the most valuable to learn experimentally given constraints on the size of the experiment and (ii) how to allocate sample size across treatment arms and sites.
@@
The first step is to specify the estimand and the underlying parameters. Define $\tau(\theta)$ the (policy-relevant) object of interest, where $\tau$ is a known smooth function and $\theta$ is a vector of unknown parameters; $\tau$ may be scalar or vector-valued. Researchers also observe external baseline estimates of $\theta$–henceforth observational estimates–which may be biased. Each feasible experiment can identify a subset of parameters without bias. This parameterization makes explicit which components of $\theta$ are learned experimentally and which may rely on (biased) observational evidence; not all components need or can be learned experimentally.
@@
Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in rural Kenya receive the subsidy. The entries of $\theta$ consist of (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium, and (iii) wage adjustments that shift the returns to schooling. Due to cost constraints, researchers may only run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$. They may choose between a CCT arm to estimate the direct effect, or an unconditional-cash arm (UCT) to measure the income elasticity, and extrapolate the GE effects using evidence from a previous study in Mexico (Todd and Wolpin, 2006). (When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ targets the average effect across sites.)
@@
The second step is to select the experiment(s), sample size and estimator. A natural starting point would be to design the experiment to minimize the mean-squared error (MSE) for $\tau(\theta)$. However, in practice, the bias is unknown prior to the experiment. A worst-case approach would be overly-conservative and disregard information about the variance. We instead adapt the definition of adaptation regret, previously studied for combining two estimators under a fixed design (Armstrong et al., 2024; Tsybakov, 1998), to our experimental design problem (with multiple parameters). Here, the regret is the ratio of the researcher’s worst-case mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the estimator. Taking its supremum over the worst-case bias yields a robust procedure without requiring a prior about the bias or on its bound. However, our task of also choosing the experimental design and the lack of an unbiased estimate for $\tau(\theta)$ when not all parameters can be learned experimentally make existing solutions not applicable.
@@
Our main result provides an explicit and novel characterization of adaptation regret for (asymptotically) linear estimators in experimental and observational moments. These include generalized methods of moments where we may also optimize over the weighting matrix. For any design and weighting matrix that combine the experimental and observational point estimates, the regret is the maximum of two normalized components. The first is a variance regret: the estimator’s sampling variance under the chosen design and weights, divided by the smallest attainable variance. The second is a bias regret: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias over the class of designs. The variance regret depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the weighting matrix, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters evaluated, under mild conditions, at the observational estimates. [^6] For this result to hold, we assume that $\tau(\theta)$ is twice differentiable with bounded derivatives. For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al. (2020); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error. Section 5.3 provides details. The bias regret depends only on the sensitivity vector and the weighting matrix, known ex-ante.
@@
Our characterization makes the bias-variance trade-off intuitive. At the optimum, the two normalized components tend to be equalized: when the bias dominates, the design prioritizes high-sensitivity parameters; when variance dominates, the procedure invests sample where it most reduces variance. This yields a nested procedure that jointly selects the estimator, the sample allocation and the experiment(s) to run, trading off precision with bias sensitivity. The resulting design and estimator only require knowledge of the expected variance-covariance matrix and observational estimates. It can be reported in a pre-analysis plan. That is, for given observational estimates, the inputs are the same as those required by standard power-analysis plans for $\tau(\theta)$ (e.g. Gerber and Green, 2012).
We then generalize our framework to general moment selection problems that combine experimentally identified moments with observational (misspecified) moments. With parameters defined by non-linear but smooth moments our results directly apply under standard local asymptotics, where the bias is local to zero at same or possibly different rate than the standard error (e.g. Andrews et al., 2017, 2020). We also generalize to partially identified $\tau$ and settings where researchers have priors (confidence sets) about the bias.
From a theoretical (and computational) perspective, our key insight is that, as we restrict the class of estimators to be (asymptotically) linear in the moments, the worst-case adaptation regret is quasi-convex in the radius of the bias, for any norm characterizing the bias’ uncertainty set. This result yields a transparent variance-bias trade-off and differs from existing solutions under the fixed-design optimal estimators as in Armstrong et al. (2024), Tsybakov (1998); there, quasi-convexity fails due to non-linearity of the estimator, and optimization corresponds to a minimax program over worst-case priors. In our joint design and estimation problem, quasi-convexity substantially simplifies optimization which is quadratic for the choice of the estimator, with a closed form expression for sample allocation and a mixed-integer formulation for the experiment choice only (typically from a finite set). This program is solved via off-the-shelf mixed-integer quadratic programming.
We illustrate the framework in a cash‐transfer application in Kenya, where the researcher combines new experimental evidence with external evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006; Attanasio et al., 2012), where these preliminary estimates may lack external validity. We compare three designs and their combinations: (i) a CCT arm to estimate direct schooling effects; (ii) a UCT arm to estimate income effects; and (iii) an employment program that identifies individual wage elasticities. Experimentation costs are binding, as cash transfers and employment programs require distinct infrastructures and implementation teams. We show that the wage experiment is the most informative for sufficiently large sample sizes, while for smaller samples a combination of CCT and UCT—allocating most of the sample to the CCT arm—is optimal. With one thousand participants, relative to the variance-minimizing (Neyman) design, our optimal two-arm design delivers more than a 400% reduction in bias (and adaptation regret) with variance no larger than $\sim$30% of the Neyman variance; for a single-arm design, our procedure yields a 140% reduction in bias at the cost of only about a 12% increase in variance. These comparisons illustrate the benefit of our method in balancing robustness and precision.
@@
As a second application, we study where to run a microfinance experiment in India, integrating evidence from earlier nonrandomized microfinance introductions. We target the average effect in the entire region, but constraint the researcher to choose one or two smaller areas for randomization due to coordination (fixed) costs constraints. We use observational estimates in Banerjee et al. (2024) to select one or more areas for randomization. Because Banerjee et al. (2024) also implement a separate randomized expansion, we can calibrate performance of our design (agnostic of the bias) by comparing each candidate design’s MSE to that of an oracle that knows the bias, using the experimental estimates to calibrate the bias. Compared to the variance-optimal benchmark, we reduce MSE by more than 250%.
@@
In summary, we complement this literature by studying which experiment to run (and how) when not all parameters are learned experimentally, a common problem in economics.
Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design. References include Manski and Tetenov (2016), Banerjee et al. (2020), Manski (2004), Dominitz and F. Manski (2017), Olea et al. (2024), Hu et al. (2024), Breza et al. (2025) among others. These references focus on settings where researchers only have access to experimental variation, instead of combining it with external (observational) evidence, motivating different designs and objective functions. We also complement literature that leverages correctly-specified models for site-selection (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
A related line of work analyzes estimation under misspecification (Armstrong et al., 2024; Armstrong and Kolesár, 2018; Andrews et al., 2025; Andrews et al., 2020; Bonhomme and Weidner, 2022; Donoho, 1994; Kwon and Sun, 2025), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024). We provide a novel characterization of adaptation regret as a quasi-convex function in the bias bound (and therefore as the maximum of the variance and bias regret) by leveraging linearity of the estimator. Such characterization shows how, in our problem, misspecification is captured through the first-order sensitivity of the estimand to each parameter, an important measure for sensitivity analysis in Andrews et al. (2020); it provides a decision-theoretic foundation for simultaneously protecting against a correctly specified and misspecified model (e.g. Bickel, 1984; Kempthorne, 1988), here through the maximum of a normalized variance and bias components. Unlike the references above, where the design is fixed and researchers optimize over the choice of the estimator, our main contribution is to optimize the design itself. This changes the optimization and regret, computed against the best design-estimator.
2 Baseline scenario: problem description
We will focus first on a simple shrinkage estimator and a univariate estimand. Our results generalize directly to general methods of moments and multivariate targets in Section 4.
Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional unknown parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$. The goal is to construct an estimator $\hat{\tau}$ that accurately approximates $\tau(\theta)$ by combining existing evidence (e.g., observational studies) with experimental variation designed by the researcher. Our question is how to design such experiments under feasibility constraints.